OpenStudy (anonymous):

simplify the expression 4x^2-12x/x^2+x-12

OpenStudy (anonymous):

is 4x^2-12x all over x^2+x-12? or is it -12x/x^2?

OpenStudy (anonymous):

its 4x^2-12x all over x^2+x-12

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

\[\left( 4x^2-12x \right)\div \left( x^2+x-12 \right)\]

OpenStudy (anonymous):

factor first.

OpenStudy (anonymous):

\[4x\left( x-3 \right)\div \left( x-3 \right)\left( x+4 \right)\]

OpenStudy (anonymous):

cancel out the (x-3)

OpenStudy (anonymous):

4x(x-3) all over x^2+x-12?

OpenStudy (anonymous):

\[4x \div(x+4)\]

OpenStudy (anonymous):

oh i can factor that

OpenStudy (anonymous):

THat's ur answer- can't do anymore from there.

OpenStudy (anonymous):

so the neg on the top and the bottom can cencel eachother out"?

OpenStudy (anonymous):

u talking bout the (x-3)?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yup. so long as the term's the same on the top and the bottom, they cancel. you knoe how 1/2=5/10, because you multiplied it by 5/5? That's pretty much the same logic- (x-3)/(x-3)=1

OpenStudy (anonymous):

an when u multiply something by 1, nothing changes